# The Signals of Doomsday I: False Higgs vacuum decay signatures

### v4

@@ 1 Introduction
It is believed that our universe has so far undergone a number of phase transitions at different energy scales, e.g. GUT, electroweak, QCD… Phase transitions usually introduce drastic changes in the structure of the universe, and if they happen late enough, they would almost certainly be fatal to any existing life forms. Until recently, late-time phase transitions have not been a topic of intensive study (see however [101, 58, 68, 5, 50, 44, 45, 16, 10, 11, 97, 51]). A similar late–time analysis of $SU(3)_{c}$ symmetry breaking in a true–vacuum background has been presented in Ref. [96] and an analysis for $U(1)_{\rm EM}$ symmetry breaking will appear soon.
After the Higgs discovery, it became apparent that a scenario similar to the one outlined in [58] (that the Higgs potential might develop instability with our current universe sitting in the unstable false vacuum) might actually be true [42, 48, 53, 99, 46]. This might not be an immediate problem, since the lifetime of our universe in the false vacuum could be many billions of years. However, as shown in [34, 33, 17, 60, 28, 29, 62, 30, 102], small primordial black holes can play the role of Higgs vacuum decay catalysers and significantly increase the tunneling probability. Depending on the small black hole mass and the exact parameters in the Higgs potential, the tunneling probability could be high enough to produce bubbles of true vacuum within the present lifetime of the universe.
@@
We investigate here the process in which the Higgs tunneling can leave a signature in the form of produced particles that can potentially reach us before the bubble. In particular, we calculate the spectrum of the Higgs particles created due to the vacuum mismatch inside and outside of the bubble. If there is no friction with the environment, this mechanism will continue indefinitely during the bubble expansion and lead to enormous particle production. Since the bubble walls expand almost with the speed of light, the signal will hit us practically at the same time as the wall itself. Therefore, to quantify the amount of particles that can reach us before the bubble wall, we introduce friction between the wall and the environment, which leads to a terminal wall velocity slower than the speed of light. We then integrate particle production up to the moment when the terminal velocity is reached.
@@ 2 The Higgs vacuum decay
The Higgs potential $V(\phi)$ from Eq. (2.2) is shown in Fig. 1, while the solution for the Higgs field distribution in the presence of gravity in Fig. 2. Since we have freedom in choosing the exact values of parameters, we follow [27], and for illustration we set $b=10^{-4}$, $\lambda=-0.001$, $\phi_{*}=0.5M_{p}$, and $\lambda_{6}=0.34$.
@@ 4 Particle production due to vacuum mismatch
The mismatch of vacua is usually accompanied by particle production. In our case, the false Higgs vacuum is outside the bubble, while the true vacuum is inside. As the bubble expands through space, the Higgs field vacuum abruptly changes from false to true. Bogoliubov transformations then indicate that this vacuum mismatch will lead to the Higgs particle production. Since the Higgs particle is very heavy in the new vacuum, its decay products will be very energetic. For simplicity, we assume that the Higgs field tunnels to the true vacuum state homogeneously. This should be a good enough approximation since particles are produced locally, i.e. the relevant phase transition length scale is smaller than the size of the bubble at virtually any time after the initial bubble nucleation. Here, for the relevant length scale we take the width of the potential barrier in Fig. 1 ($\sim[10^{-2}M_{p}]^{-1}$), which is also the scale of the Higgs expectation value in the true vacuum.
@@
$\partial_{\tilde{\tau}}^{2}h+\nabla^{2}h-\mu^{2}h=0\text{, for $\tilde{\tau}>\tilde{\tau}^{*}$}$, (4.3)
@@ 9 Conclusions
We note that the crucial assumption here is that there is no new physics beyond the standard model. New physics could potentially modify the Higgs potential and render our vacuum stable. Another possibility is that small primordial black holes do not exist. Since they are necessary catalysers, their absence in an appropriate mass range [36] would perhaps invalidate this doomsday scenario. We also stress that even if Higgs-instability regions are seeded in the Universe, their fate need not be “doomsday” in all circumstances: De Luca–Kehagias–Riotto [41] argue that in the thin-wall regime such dangerous patches can collapse due to Higgs backreaction (often leaving black holes), rather than expanding and eating the Universe, which can substantially relax the corresponding inflationary Hubble bound.
@@ Appendix A Appendix A: Proper Acceleration of the Bubble Wall
Thus, in the absence of friction, the magnitude of the proper acceleration of the bubble wall is constant and given by

### v5

@@ 1 Introduction
It is believed that our universe has so far undergone a number of phase transitions at different energy scales, e.g. GUT, electroweak, QCD… Phase transitions usually introduce drastic changes in the structure of the universe, and if they happen late enough, they would almost certainly be fatal to any existing life forms. Until recently, late-time phase transitions have not been a topic of intensive study (see however [112, 61, 72, 6, 53, 47, 48, 17, 11, 12, 108, 54, 91, 45, 90, 89, 44, 74, 73, 86, 79, 75, 2]). A similar late–time analysis of $SU(3)_{c}$ symmetry breaking in a true–vacuum background has been presented in Ref. [107] and an analysis for $U(1)_{\rm EM}$ symmetry breaking will appear soon.
After the Higgs discovery, it became apparent that a scenario similar to the one outlined in [61] (that the Higgs potential might develop instability with our current universe sitting in the unstable false vacuum) might actually be true [43, 51, 56, 110, 49, 68]. This might not be an immediate problem, since the lifetime of our universe in the false vacuum could be many billions of years. However, as shown in [35, 34, 18, 63, 29, 30, 65, 31, 113], small primordial black holes can play the role of Higgs vacuum decay catalysers and significantly increase the tunneling probability. Depending on the small black hole mass and the exact parameters in the Higgs potential, the tunneling probability could be high enough to produce bubbles of true vacuum within the present lifetime of the universe. In this context, one should distinguish between the scale at which the Higgs quartic first turns negative and the generally higher field value at which a stabilized true minimum may appear once higher-dimensional operators are included; the latter is the relevant scale for the benchmark considered in this work.
@@
We investigate here the process in which the Higgs tunneling can leave a signature in the form of produced particles that can potentially reach us before the bubble. In particular, we calculate the spectrum of the Higgs particles created due to the vacuum mismatch inside and outside of the bubble. A natural question is why we focus here on Higgs production, rather than on other possible channels such as direct production of lighter fermion pairs or interactions of ambient particles in the interstellar or intergalactic medium with the bubble wall. Our main reason is that the transition under consideration is itself a Higgs-field transition, so Higgs excitations provide the most direct probe of the changing vacuum background. In that sense, the Higgs sector is the most immediate channel to analyze in a first phenomenological study of false-vacuum decay signatures.
A second reason is that Higgs production provides a comparatively direct bridge between the microscopic bubble dynamics and the long-range messengers relevant for observation. Once produced, the heavy Higgs excitations promptly decay into Standard Model particles, whose subsequent decay chains, hadronization, annihilation, and cascades generate photons and neutrinos. These are precisely the observables of interest in the present work.
We do not assume that Higgs bosons are universally the dominant channel in every physical environment. Other processes, including direct production of additional particle species from the time-dependent vacuum background and interactions of the bubble wall with ambient matter, may also contribute and in some settings could be important. However, a systematic treatment of all such channels would substantially broaden the scope of this paper. Our aim here is therefore more limited: to isolate and study one particularly direct and physically transparent channel associated with the Higgs-sector transition itself. For this reason, Higgs production serves as the natural starting point for the present analysis, while a more complete comparison with other production mechanisms is left for future work.
If there is no friction with the environment, this mechanism will continue indefinitely during the bubble expansion and lead to enormous particle production. Since the bubble walls expand almost with the speed of light, the signal will hit us practically at the same time as the wall itself. Therefore, to quantify the amount of particles that can reach us before the bubble wall, we introduce friction between the wall and the environment, which leads to a terminal wall velocity slower than the speed of light. We then integrate particle production up to the moment when the terminal velocity is reached.
@@ 2 The Higgs vacuum decay
The Higgs potential $V(\phi)$ from Eq. (2.2) is shown in Fig. 1, while the solution for the Higgs field distribution in the presence of gravity in Fig. 2. Since we have freedom in choosing the exact values of parameters, we follow [28], and for illustration we set $b=10^{-4}$, $\lambda=-0.001$, $\phi_{*}=0.5M_{p}$, and $\lambda_{6}=0.34$. It is important to distinguish between two different field scales that appear in discussions of Higgs vacuum metastability. The scale often quoted in the literature, of order $10^{10}\,\mathrm{GeV}$ [54], refers to the field value at which the renormalization-group improved Higgs quartic coupling first turns negative, signalling the onset of metastability. In the present work, however, the physically relevant scale is not that crossing point itself, but the location of the stabilized high-field true minimum of the effective potential. Because we include the higher-dimensional operator in Eq. (2.2), the potential develops a true minimum at a substantially larger field value. It is this larger field value that controls the Higgs mass in the bubble interior, the wall thickness, and the characteristic energies of the particles emitted from the bubble wall.
Our benchmark choice is also tied to the black-hole-catalysed vacuum decay scenario. In that setup, the bubble nucleation process is seeded by a small primordial black hole, and the stabilized true minimum provides the relevant scale entering the bounce configuration and the subsequent bubble dynamics. We therefore emphasize that the lower instability scale marks the onset of Higgs metastability, whereas the larger fiducial field value used throughout this paper corresponds to the location of the stabilized true vacuum employed in our benchmark analysis.
@@ 4 Particle production due to vacuum mismatch
The mismatch of vacua is usually accompanied by particle production. In the present work we focus on the Higgs channel because it is the field directly undergoing the vacuum transition, and therefore provides the most immediate probe of the changing vacuum configuration. In our case, the false Higgs vacuum is outside the bubble, while the true vacuum is inside. As the bubble expands through space, the Higgs field vacuum abruptly changes from false to true. Bogoliubov transformations then indicate that this vacuum mismatch will lead to the Higgs particle production. Since the Higgs particle is very heavy in the new vacuum, its decay products will be very energetic. For simplicity, we assume that the Higgs field tunnels to the true vacuum state homogeneously. This should be a good enough approximation since particles are produced locally, i.e. the relevant phase transition length scale is smaller than the size of the bubble at virtually any time after the initial bubble nucleation. Here, for the relevant length scale we take the width of the potential barrier in Fig. 1 ($\sim[10^{-2}M_{p}]^{-1}$), which is also the scale of the Higgs expectation value in the true vacuum.
@@
$\partial_{\tilde{\tau}}^{2}h+\nabla^{2}h-\mu^{2}h=0\text{, for $\tilde{\tau}>\tilde{\tau}^{*}$}$,
@@ 7 Thermal particle production from frictional dissipation
In principle, once the wall enters a quasi-steady-state regime, direct scattering and momentum transfer of ambient particles across the moving wall can provide an additional source of particle production, as emphasized in microscopic studies of bubble-wall friction and transition-radiation effects. However, in the present setup we expect this contribution to be subdominant compared with the thermal particle production generated by the much larger frictional energy dissipated into the shocked medium, and we therefore do not consider it here.
@@ 9 Conclusions
We note that the crucial assumption here is that there is no new physics beyond the standard model. New physics could potentially modify the Higgs potential and render our vacuum stable. Another possibility is that small primordial black holes do not exist. Since they are necessary catalysers, their absence in an appropriate mass range [37] would perhaps invalidate this doomsday scenario. We also note that our quantitative benchmark depends on the location of the stabilized true minimum of the Higgs potential. If the true-vacuum scale were lower than in the benchmark adopted here, then all characteristic mass scales in the problem would be correspondingly reduced. In particular, the Higgs mass in the true vacuum would be smaller, the bubble wall thickness would typically be larger, and the resulting photon and neutrino spectra would be shifted toward lower characteristic energies. The qualitative picture developed in this work would remain unchanged — namely, bubble expansion, vacuum-mismatch particle production, friction-induced thermal production, and long-range photon/neutrino signatures — but the signal would be softer and the detailed phenomenology would have to be re-evaluated for that lower-scale vacuum structure. We also stress that even if Higgs-instability regions are seeded in the Universe, their fate need not be “doomsday” in all circumstances: De Luca–Kehagias–Riotto [42] argue that in the thin-wall regime such dangerous patches can collapse due to Higgs backreaction (often leaving black holes), rather than expanding and eating the Universe, which can substantially relax the corresponding inflationary Hubble bound.
@@ Appendix A Appendix A: Proper Acceleration of the Bubble Wall
Thus, in the absence of friction, the magnitude of the proper acceleration of the bubble wall is constant and given by [59, 118]